# Factor Risk Budgeting and Beyond

### v1

@@ 3.1 Linear factor models and factor exposures
It is common practice for portfolio and risk managers to write $X$ as a linear function of $m$ factor returns modelled by $F\in L^{0}(\Omega,\mathbb{R}^{m})$ and $d$ idiosyncratic components modelled by $\epsilon\in L^{0}(\Omega,\mathbb{R}^{d})$ where $F$ and $\epsilon$ are assumed to be independent. This modelling framework is commonly known as a linear factor model and is mathematically expressed as:
$X=\beta F+\epsilon$ (3)
@@
Factor loadings characterize the sensitivity of assets to specific risk factors. They provide a quantitative measure of how much the return of an asset are anticipated to fluctuate in response to a one-unit change in a particular risk factor. These risk factors encompass a broad spectrum, including economic factors such as economic growth and inflation, stock-specific features like size, value, quality and profitability, and statistical factors extracted from data. Factor loadings are typically estimated through statistical techniques like regression analysis. They can also be estimated based on sophisticated proprietary quantitative models.
@@
If we consider a model as in Eq. (3), then the portfolio loss, for a given vector of asset exposures $y$, can be written as
$-y^{\prime}X=-w^{\prime}F-y^{\prime}\epsilon$
@@
Under such a model, one is mostly interested in managing the portfolio risk through the factor exposures and achieving diversification among factors rather than solely focusing on the exposures of the individual assets since their returns are actually dependent on the returns of the underlying factors. Therefore, it makes complete sense to try to achieve a desired level of factor exposures $w$ by finding a vector of asset exposures $y$ that satisfies $w=\beta^{\prime}y$. However, an issue arises due to the existence of an infinite number of asset exposure vectors that meet this condition.
One of these infinitely many vectors is of course $\bar{y}_{w}=\beta(\beta^{\prime}\beta)^{-1}w$ [^6] For instance, [29] considers this specific case. Therefore, we clearly see that $\beta^{\prime}y=w\iff\beta^{\prime}y=\beta^{\prime}\bar{y}_{w}\iff y-\bar{y}_{w}\in\text{Ker}(\beta^{\prime})=\text{Im}(\beta)^{\perp}$. Thus, the set of asset exposures with the factor exposures $w$ is the affine subspace $\bar{y}_{w}+\text{Im}(\beta)^{\perp}$. Amongst all the vectors $y\in\bar{y}_{w}+\text{Im}(\beta)^{\perp}$ of asset exposures compatible with a vector of factor exposures $w$, some may be more favourable than others when considering additional criteria that are important for portfolio managers. In the following section, we address this concern and introduce a risk-based framework to tackle this issue.
@@ 3.2 From factor exposures to portfolio risk: a factor risk measure
Given a full-rank matrix $\beta\in\mathbb{R}^{d\times m}$ and an asset exposure risk measure $\mathcal{R}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying the assumptions of Section 2, we can define the factor risk measure $\mathcal{S}_{\mathcal{R},\beta}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ (hereafter simply denoted by $\mathcal{S}$) by
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where $m=\underset{x\in\mathbb{R}^{d},\lVert x\rVert=1}{\min}\mathcal{R}(x)>0.$ So, $\forall p\in\bar{B}(0,m)$, *$\mathcal{R}^{*}(p)=0.$
@@
(*$\partial(\mathcal{R}^{*}+\mathcal{X}_{\text{Im}(\beta)})(p)=\partial\mathcal{R}^{*}(p)+\partial\mathcal{X}_{\text{Im}(\beta)}(p)=\partial\mathcal{R}^{*}(p)+\text{Im}(\beta)^{\perp}.$
@@
Given a vector $w$ of factor exposures, Euler homogeneous function theorem leads to a breakdown of risk as999 As stated in Proposition 4, the risk contribution of the factor $i$, i.e. $w_{i}\partial_{i}\mathcal{S}(w)$, can be easily calculated using $\nabla\mathcal{R}$. When the risk measure is volatility, this gradient has an analytic form. As for the sensitivity of Expected Shortfall, general expressions exist (see [15] and [33]).
@@ 4.1 Framing the problem of Factor Risk Budgeting
By analogy with Risk Budgeting, we can use $\mathcal{S}$ to define a Factor Risk Budgeting portfolio allocating risk among the underlying risk factors given factor risk budgets. Mathematically, we define the Factor Risk Budgeting (FRB) problem as follows:
@@
[…]
@@
Factor Risk Budgeting portfolios are not necessarily long-only portfolios.
@@
[…]
@@ 4.2 The case of long-only portfolios
It is essential to note that Factor Risk Budgeting portfolios presented above constitute fully invested general portfolios. In practice, short positions are not always permitted, and asset weights generally have to comply with certain constraints. Especially, it is often the case that portfolios are restricted to be long-only. We therefore highlight two possible approaches for the case of long-only portfolios in an FRB setting.
Solving Problem (4.1) or (5) results in an FRB portfolio $\theta\in\Delta_{d}$ and a vector of factor exposures $w^{*}$ that satisfies the condition set out in the Definition 3. Since $\theta$ is not necessarily long-only, one possible approach is to look for an alternative portfolio that respects the constraints and produces the same factor exposures $w^{*}$. Precisely, this can be done on the basis of our initial risk-based approach by solving
@@
Another approach can therefore be to incorporate the long-only constraint directly in Problem (4.1) or (5) and consider the set $\mathcal{C}^{\geq 0}=\big{\{}y\in\mathbb{R}_{+}^{d}|(\beta^{\prime}y)_{i}>0,\forall i\in\{1,\ldots,m\}\big{\}}$ for the variable $y$ [^10] Even the set $\mathcal{C}^{\geq 0}$ is not always non-empty, meaning that there may not even exist a nonnegative vector of asset exposures leading to a positive vector of factor exposures. A simple approach to establish this by contradiction is to consider $\beta$ whose elements are negative. In such a scenario, no nonnegative vector $y$ can lead to a positive vector $w=\beta^{\prime}y$. However, this constraint is more of a theoretical limitation than a practical one: in most practical situations, $\mathcal{C}^{\geq 0}$ is a non-empty set. There, of course, still exists a unique minimizer that solves the aforementioned problem. Nevertheless, $\theta:=y^{*}/\sum_{i=1}^{d}y^{*}_{i}\in\Delta^{\geq 0}_{d}$, now, may not guarantee a solution to $\text{FRB}_{b}$. On the other hand, we expect this to be a portfolio in which the factor risk contributions are close to the factor risk budgets. We will examine this approach in the numerical section and analyse the proximity of factor risk contributions to factor risk budgets.
4.3 Managing the interplay: balancing asset and factor risk contributions
Although FRB portfolios may present a theoretical appeal, they can run into a number of practical pitfalls. The search for a precise match in factor risk budgets can lead to concentrated portfolios in asset weights, especially in the original long-short setting, amplifying individual asset risk contributions. Likewise, classical RB portfolios may contain hidden risks associated with the underlying risk factors since the sole focus is to satisfy desired risk contribution from the assets.
Consequently, a framework benefiting from both ideas can be of practical use to construct portfolios which take asset and factor risk contributions into account together in the same setting. The fact that FRB portfolios can be found by solving an optimization problem involving only $\mathcal{R}$ allows us to propose an optimization framework that seeks a balanced solution between FRB and conventional RB portfolios.
@@ 4.3 Managing the interplay: balancing asset and factor risk contributions
$\min_{y\in\mathcal{C}^{>0}}g(\mathcal{R}(y))-\lambda_{a}\sum_{i=1}^{d}{b_{a}}_{i}\log{y_{i}}-\lambda_{f}\sum_{i=1}^{m}{b_{f}}_{i}\log{(\beta^{\prime}y)_{i}}$ (6)
where $\mathcal{C}^{>0}=\big{\{}y\in(\mathbb{R}_{+}^{*})^{d}|(\beta^{\prime}y)_{i}>0,\forall i\in\{1,\ldots,m\}\big{\}}$ and $\lambda_{a},\lambda_{f}\in\mathbb{R}_{+}^{*}$ are asset and factor importance parameters. One can introduce different levels of sensitivities to the different types of risk contributions by choosing importance parameters accordingly.
@@
If $(y^{*},\zeta^{*})$ is a solution of the above stochastic optimization problem, then, $\theta:=y^{*}/\sum_{i=1}^{d}y^{*}_{i}$ is called the Asset-Factor Risk Budgeting portfolio.
@@
We should note that Asset-Factor Risk Budgeting portfolios never exactly match given risk budgets for assets and factors. However, this formulation allows us to explore a compromise solution that reconciles the advantages of FRB portfolios as well as penalizing large deviations from asset risk budgets.
5 Numerical results
5.1 Analysis of risk-based portfolios under an equity factor model

### v2

@@ 3.1 Linear factor models and factor exposures
It is common practice for portfolio and risk managers to write $X$ as a linear function of $m$ factor returns modelled by $F\in L^{0}(\Omega,\mathbb{R}^{m})$ and $d$ idiosyncratic components modelled by $\epsilon\in L^{0}(\Omega,\mathbb{R}^{d})$ where $F$ and $\epsilon$ are assumed to be independent. This modelling framework is commonly known as a linear factor model and is mathematically expressed as
$X=\beta F+\epsilon,$ (3)
@@
Factor loadings characterize the sensitivity of assets to specific risk factors. They provide a quantitative measure of how much the return of an asset are expected to fluctuate in response to a one-unit change in a particular risk factor. These risk factors encompass a broad spectrum, including economic factors such as economic growth and inflation, stock-specific features like size, value, quality and profitability, and statistical factors extracted from data. Factor loadings are typically estimated through statistical techniques like regression analysis.
@@
In the modeling framework of Eq. (3), the portfolio loss for a given vector of asset exposures $y$ is
$-y^{\prime}X=-w^{\prime}F-y^{\prime}\epsilon,$
@@
In such a model, one is mostly interested in managing the portfolio risk through factor exposures and achieving diversification among factors rather than solely focusing on the exposures to the individual assets since their returns are actually dependent on the returns of the underlying factors. Therefore, it makes sense to try to achieve a desired level of factor exposures $w$ by finding a vector of asset exposures $y$ that satisfies $w=\beta^{\prime}y$. However, an issue arises due to the existence of an infinite number of asset exposure vectors that meet this condition.
Because $\beta$ is full rank, one of these infinitely many vectors is $\bar{y}_{w}=\beta(\beta^{\prime}\beta)^{-1}w$ [^9] For instance, [33] considers this specific case. Therefore, we clearly see that $\beta^{\prime}y=w\iff\beta^{\prime}y=\beta^{\prime}\bar{y}_{w}\iff y-\bar{y}_{w}\in\text{Ker}(\beta^{\prime})=\text{Im}(\beta)^{\perp}$. Thus, the set of asset exposures with the factor exposures $w$ is the affine subspace $\bar{y}_{w}+\text{Im}(\beta)^{\perp}$. Amongst all the vectors $y\in\bar{y}_{w}+\text{Im}(\beta)^{\perp}$ of asset exposures compatible with a vector of factor exposures $w$, some may be more favourable than others when considering additional criteria that are important for portfolio managers. In the following section, we address this concern and introduce a risk-based framework to tackle this issue.
@@ 3.2 From factor exposures to portfolio risk: a factor risk measure
For a full-rank matrix $\beta\in\mathbb{R}^{d\times m}$ and an asset exposure risk measure $\mathcal{R}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying the assumptions of Section 2, we define the factor risk measure $\mathcal{S}_{\mathcal{R},\beta}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ (hereafter simply denoted by $\mathcal{S}$) by
$\mathcal{S}:w\mapsto\inf_{y\in\bar{y}_{w}+\text{Im}(\beta)^{\perp}}\mathcal{R}(y)=\inf_{u\in\text{Im}(\beta)^{\perp}}\mathcal{R}(\beta(\beta^{\prime}\beta)^{-1}w+u).$
@@
$\mathcal{S}$ is convex as it can be expressed with the infimal convolution [^10] We recall that the infimal convolution of two functions $\phi$ and $\psi$ is defined as $\phi\square\psi(a)=\inf\{\phi(a-b)+\psi(b)|b\in\mathbb{R}^{d}\}.$ of $\mathcal{R}$ and $\mathcal{X}_{\text{Im}(\beta)^{\perp}}$:
$\mathcal{S}(w)=\mathcal{R}\square\mathcal{X}_{\text{Im}(\beta)^{\perp}}(\beta(\beta^{\prime}\beta)^{-1}w),$
@@
$\forall y\in\mathbb{R}^{d}$, there exists a unique $u^{*}(y)\in\text{Im}(\beta)^{\perp}$ such that
$\inf_{u\in\text{Im}(\beta)^{\perp}}\mathcal{R}(y+u)=\mathcal{R}(y+u^{*}(y)).$
@@
The convex function $\mathcal{R}\square\mathcal{X}_{\text{Im}(\beta)^{\perp}}$ is proper, positively homogeneous on $\mathbb{R}^{d}$ and continuously differentiable on $\mathbb{R}^{d}\setminus\text{Im}(\beta)^{\perp}$ with
$\forall y\in\mathbb{R}^{d}\setminus\text{Im}(\beta)^{\perp},\quad\nabla(\mathcal{R}\square\mathcal{X}_{\text{Im}(\beta)^{\perp}})(y)=\nabla\mathcal{R}(y-u^{*}(y)).$
@@
where $m=\underset{x\in\mathbb{R}^{d},\lVert x\rVert=1}{\min}\mathcal{R}(x)>0.$ So, $\forall p\in\bar{B}(0,m)$, $\mathcal{R}^{*}(p)=0.$
@@
$\partial(\mathcal{R}^{*}+\mathcal{X}_{\text{Im}(\beta)})(p)=\partial\mathcal{R}^{*}(p)+\partial\mathcal{X}_{\text{Im}(\beta)}(p)=\partial\mathcal{R}^{*}(p)+\text{Im}(\beta)^{\perp}.$
@@
Given a vector $w$ of factor exposures, Euler homogeneous function theorem leads to a breakdown of risk as121212 As stated in Proposition 4, the risk contribution of the factor $i$, i.e. $w_{i}\partial_{i}\mathcal{S}(w)$, can be easily calculated using $\nabla\mathcal{R}$. When the risk measure is volatility, this gradient has an analytic form. As for the sensitivity of Expected Shortfall, general expressions exist (see [18] and [37]).
@@ 4.1 Framing the problem of Factor Risk Budgeting
By analogy with Risk Budgeting, we can use $\mathcal{S}$ to define a Factor Risk Budgeting portfolio allocating risk across the underlying risk factors given factor risk budgets. Mathematically, we define the Factor Risk Budgeting (FRB) problem as follows:
@@
Let $\beta\in\mathbb{R}^{d\times m}$ be a full rank matrix of factor loadings. Let $b\in\Delta^{>0}_{m}$ be a vector of factor risk budgets. We say that a vector of weights $\theta\in\Delta_{d}$, solves the Factor Risk Budgeting problem $\text{FRB}_{b}$ if and only if
@@
In the above definition, the vector $w$ of factor exposures has positive coordinates. However, Factor Risk Budgeting portfolios are not necessarily long-only portfolios.
@@
Then, $\forall y\in y^{*}(w^{*})+\text{Im}(\beta)^{\perp}$, $\sum_{i=1}^{d}y_{i}\neq 0\implies\theta(y)=\frac{y}{\sum_{i=1}^{d}y_{i}}\in\Delta_{d}$ is solution of $\text{FRB}_{b}$.
Moreover, if $\theta\in\Delta_{d}$ is solution of $\text{FRB}_{b}$, then $w=\beta^{\prime}\theta$ is proportional to $w^{*}$.
@@ 4.2 The case of long-only portfolios
It is essential to note that Factor Risk Budgeting portfolios presented above constitute fully invested general portfolios. In practice, short positions are not always permitted, and asset weights generally have to comply with certain constraints. Especially, it is often the case that portfolios are restricted to be long only. We therefore highlight two possible approaches for the case of long-only portfolios in an FRB setting.
Solving Problem (4.1) or (5) results in an FRB portfolio $\theta\in\Delta_{d}$ and a vector of factor exposures $w^{*}$ that satisfy the condition set out in the Definition 3. Since $\theta$ is not necessarily long only, one possible approach is to look for an alternative portfolio that respects the constraints and produces the same factor exposures $w^{*}$. Precisely, this can be done on the basis of our initial risk-based approach by solving
@@
Another approach can therefore be to incorporate the long-only constraint directly in Problem (4.1) or (5) and consider the set $\mathcal{C}^{\geq 0}=\big{\{}y\in\mathbb{R}_{+}^{d}|(\beta^{\prime}y)_{i}>0,\forall i\in\{1,\ldots,m\}\big{\}}$ for the variable $y$ [^13] Even the set $\mathcal{C}^{\geq 0}$ is not always non-empty, meaning that there may not even exist a nonnegative vector of asset exposures leading to a positive vector of factor exposures. A simple approach to establish this by contradiction is to consider $\beta$ whose elements are negative. In such a scenario, no nonnegative vector $y$ can lead to a positive vector $w=\beta^{\prime}y$. However, this constraint is more of a theoretical limitation than a practical one: in most practical situations, $\mathcal{C}^{\geq 0}$ is a non-empty set. There, of course, still exists a unique minimizer that solves the aforementioned problem. Nevertheless, $\theta:=y^{*}/\sum_{i=1}^{d}y^{*}_{i}\in\Delta^{\geq 0}_{d}$, now, may not guarantee a solution to $\text{FRB}_{b}$. On the other hand, we expect this to be a portfolio in which the factor risk contributions are close to the factor risk budgets. We consider this latter approach in the numerical section of this paper and analyse the proximity of factor risk contributions to factor risk budgets for long-only FRB portfolios.
5 Balancing asset and factor risk contributions
5.1 Asset-Factor Risk Budgeting portfolios
Although FRB portfolios may present a theoretical appeal, they can run into a number of practical pitfalls. The search for a precise match in factor risk budgets can lead to concentrated portfolios in asset weights, especially in the original long-short setting, amplifying individual asset risk contributions. Likewise, classical RB portfolios may contain hidden risks associated with the underlying risk factors since the sole focus is to satisfy desired risk contribution at the asset level.
Consequently, a framework benefiting from both ideas can be of practical use to construct portfolios which take asset and factor risk contributions into account together in the same setting. The fact that FRB portfolios can be found by solving an optimization problem (Problem (4.1) or (5)) involving only $\mathcal{R}$ allows us to propose an optimization framework that seeks a balanced solution between FRB and classical RB portfolios.
@@ 5.1 Asset-Factor Risk Budgeting portfolios
$\min_{y\in\mathcal{C}^{>0}}g(\mathcal{R}(y))-\lambda_{a}\sum_{i=1}^{d}{b_{a}}_{i}\log{y_{i}}-\lambda_{f}\sum_{i=1}^{m}{b_{f}}_{i}\log{(\beta^{\prime}y)_{i}},$ (6)
where $\mathcal{C}^{>0}=\big{\{}y\in(\mathbb{R}_{+}^{*})^{d}|(\beta^{\prime}y)_{i}>0,\forall i\in\{1,\ldots,m\}\big{\}}$ and $\lambda_{a},\lambda_{f}\in\mathbb{R}_{+}^{*}$ are asset and factor importance parameters.
@@
If $(y^{*},\zeta^{*})$ is a solution of the above stochastic optimization problem, then, $\theta:=y^{*}/\sum_{i=1}^{d}y^{*}_{i}$ is called the Asset-Factor Risk Budgeting (AFRB) portfolio.
As a result, the AFRB portfolio is obtained by solving an optimization problem that involves minimizing a risk term subject to two logarithmic barriers. The risk measure and the barriers naturally prevent asset and factor exposures from diverging and approaching zero, ensuring that the resulting portfolios are balanced in terms of both exposure types and associated risks. By adjusting the importance parameters, one can introduce varying degrees of sensitivity to different types of risk contributions, thus enabling the introduction of tilts towards certain aspects.
@@
We should note that AFRB portfolios never exactly match the given risk budgets for assets and factors. However, this formulation allows us to explore a compromise solution that reconciles the advantages of FRB portfolios as well as penalizing large deviations from asset risk budgets.
5.2 Illustration with a simple model
Let us examine a simple model with 4 assets and 3 factors to gain insight into the proposed Asset-Factor Risk Budgeting framework and its distinctions from FRB and classical RB. For simplicity, we adopt the volatility risk measure and assume the following arbitrary parameters. [^14] The parameters are taken from the example presented in [33].
$\beta\!=\!\begin{bmatrix}0.9&0.0&0.5\\ 1.1&0.5&0.0\\ 1.2&0.3&0.2\\ 0.8&0.1&0.7\\ \end{bmatrix}$ and $\Sigma\!=\!\begin{bmatrix}0.0449&0.0396&0.0442&0.0323\\ 0.0396&0.0734&0.0543&0.0357\\ 0.0442&0.0543&0.0689&0.0401\\ 0.0323&0.0357&0.0401&0.0531\\ \end{bmatrix}$.
With these specifications, we can compute the RB, FRB and AFRB portfolios through Problems (1), (4.1) and (6), respectively, based on the risk budgets $b_{a}=(0.25,0.25,0.25,0.25)$ for assets and $b_{f}=(1/3,1/3,1/3)$ for factors. For AFRB, we also require the importance parameters $(\lambda_{a},\lambda_{f})$, which we set to $(0.2,0.8)$ in this simple example.
Table 1 illustrates the portfolios computed under the aforementioned setting. As expected, the RB portfolio evenly distributes its risk across the assets. In contrast, the FRB portfolio appears concentrated on two assets, both in terms of weights and risk contributions. In particular, it has the highest total risk across our three portfolios. We clearly see that the AFRB portfolio prevents excessive asset risk concentration within the portfolio.
$\theta_{i}$ $\theta_{i}\partial_{i}\mathcal{R}(\theta)$
Asset RB FRB AFRB RB FRB AFRB
1 27.86 -6.60 18.26 5.28 -1.05 3.33
2 22.60 34.95 25.72 5.28 7.93 6.00
3 21.98 8.87 17.97 5.28 1.90 4.22
4 27.56 62.78 38.05 5.28 13.38 7.63
$\sum_{1\leq i\leq 4}$ 100.00 100.00 100.00 21.13 22.16 21.18
Table 1: Asset weights and associated asset risk contributions (in %) of the portfolios computed using the different methods.
To conduct a similar analysis from the perspective of the underlying risk factors, we refer to Table 2. Firstly, for the three portfolios, we observe significant differences in risk allocation across factors, although they have similar factor exposures. Factor risk contributions in the FRB portfolio are perfectly in line with the chosen risk budgets, potentially leading to the previously observed unbalanced asset weights. Conversely, the RB portfolio, lacking sensitivity to factors by definition, is highly exposed to the risk of the first factor. The AFRB portfolio, designed with sensitivity to both assets and factors, mitigates that effect with only a negligible increase in the total portfolio risk.
$w_{i}$ $w_{i}\partial_{i}\mathcal{S}(w)$
Factor RB FRB AFRB RB FRB AFRB
1 98.36 93.37 96.73 16.64 7.39 13.87
2 20.65 26.42 22.06 1.80 7.39 3.22
3 37.62 42.42 39.36 2.66 7.39 4.08
$\sum_{1\leq i\leq 3}$ 156.63 162.21 158.15 21.11 22.16 21.17
Table 2: Factor exposures and associated factor risk contributions (in %) of the portfolios computed using the different methods.
5.3 Choosing the importance parameters
The selection of the importance parameters is a critical aspect. Ultimately, Asset-Factor Risk Budgeting aims to identify portfolios whose normalized risk contributions closely align with a reference vector of risk budgets. Hence, it is essential to assess how far we deviate from the risk budgets in terms of risk contributions for given values of $(\lambda_{a},\lambda_{f})$.
Relative entropy serves as a useful tool for conducting such analyses since risk budgets are strictly positive and sum up to 1, making them a suitable reference distribution for risk contributions. [^15] We assume nonnegative risk contributions, which is typically the case for long-only portfolios in assets and factors. In the event of negative risk contributions, one can utilize a norm as a distance measure to conduct a similar analysis. Specifically, we can compute an asset-based relative entropy score $\sum_{i=1}^{d}q_{i}\log{(q_{i}/b_{a_{i}})}$ and a factor-based relative entropy score $\sum_{i=1}^{m}s_{i}\log{(s_{i}/b_{f_{i}})}$, respectively, where $q_{i}=\theta_{i}\partial_{i}\mathcal{R}(\theta)/\mathcal{R}(\theta)$ and $s_{i}=w_{i}\partial_{i}\mathcal{S}(w)/\mathcal{S}(w)$. These scores provide insights into the degree of deviation from the target risk contributions and can be helpful to determine the importance parameters. By selecting a pair where both relative entropy scores are low, one can identify the importance parameters that result in a small amount of deviation from the desired risk allocations.
Figure 1: The impact of the choice of $(\lambda_{a},\lambda_{f})$ on the distance of normalized risk contributions from specified risk budgets at the asset level (left) and factor level (right).
Figure 1 illustrates the relative entropy scores of risk contributions for the AFRB portfolios computed across a range of different parameter values in the above model. As expected, as $\lambda_{a}$ increases, normalized asset risk contributions tend to converge towards $b_{a}$, evidenced by the asset-based relative entropy score approaching zero. However, this convergence is associated with an increase in factor-based relative entropy, indicating a shift in factor risk contributions. Conversely, to achieve normalized factor risk contributions that closely match $b_{f}$, one should choose $\lambda_{f}$ to be greater than $\lambda_{a}$. Therefore, it is crucial to ensure that the selected parameters maintain a good balance between both relative entropy scores. Additionally, one can adopt a systematic approach to conduct a grid search for a point that minimizes a combined score, e.g. the mean of asset-based and factor-based relative entropy scores.
6 Numerical applications with financial data
6.1 Analysis of risk-based portfolios in an equity factor model